OpenStudy (anonymous):

Factor: 9w^2-30w-48=0

OpenStudy (anonymous):

\[\left\{w\to \frac{1}{3} \left(5-\sqrt{73}\right),w\to \frac{1}{3} \left(5+\sqrt{73}\right)\right\} \]